# Busemann Functions in the Wasserstein Space: Existence, Closed-Forms, and Applications to Slicing

### v1

@@ 4.3 Slicing Datasets with Busemann
To leverage the closed-form between Gaussian (19), we use a Gaussian approximation of the classes, with a possible dimension reduction of the features beforehand as in (Hua et al., 2023). Let us denote $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ a dimension reduction operator with $d^{\prime}\ll d$, and, for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d^{\prime}})$, the Gaussian approximation $\Xi(\mu)=\mathcal{N}\big(m(\mu),\Sigma(\mu)\big)$ with $m(\mu)=\int x\mathrm{d}\mu(x)$ and $\Sigma(\mu)=\int(x-m(\mu))\otimes(x-m(\mu))\ \mathrm{d}\mu(x)$ the mean and covariance operators. The label projections are then given by $Q^{\eta}(y)=B^{\eta}\big(\Xi\big(\psi_{\#}\varphi(y)\big)\big)$ with $\eta$ a geodesic ray on $\mathrm{BW}(\mathbb{R}^{d})$.
@@
with $\phi$ the cumulative distribution function of $\mathcal{N}(0,1)$. In practice, the geodesic ray can be sampled using $m_{1}\sim\mathcal{U}([-1,1])$ and setting $\sigma_{1}=\sqrt{1-m_{1}}$. Given $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, we define the Sliced-Wasserstein Busemann 1D Gaussian distance ($\mathrm{SWB1DG}$) as
@@ 5 EXPERIMENTS
5.1 Correlation with OTDD
We report the results in Figure 1, where we first scatter the values obtained in ordinate for OTDD, and in abscissa for the sliced distances with 5000 projections. Then, we report the values of the Spearman and Pearson correlations.The Pearson correlation is equal to $\pm 1$ when both quantities are linearly correlated while Spearman correlation is equal to $\pm 1$ if the quantities are monotonically related. We observe that both $\mathrm{SWB1DG}$ and $\mathrm{SWBG}$ are better correlated to OTDD than $\mathrm{SOTDD}$. We note that the results do not match those in (Nguyen et al., 2025a), where they used only 10 pairs. We hypothesize that for such small number of samples, the Spearman and Pearson correlation are very sensitive to randomness. To verify the robustness of our results, we report in Table 1 the results obtained by bootstrapping 50 pairs of subdatasets, and averaging over 10 experiments. Using this process for different numbers of projections, we see that the Busemann based distances always outperform $\mathrm{SOTDD}$, and often under a much smaller number of projections.
@@ 5.1 Correlation with OTDD
Figure 2: Evolution of the WoW gradient flow of SWBG and SOTDD with the three rings dataset as target.
Figure 3: Convergence of the flow towards the 3 rings, averaged over 100 random batches of the target.
@@ 5.2 Flowing Labeled Datasets
Bonet et al. (2025a) modeled this task as a minimization problem over the space $\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, representing datasets with $n$ samples by class as $\mathbb{P}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ with $\mu_{c}=\frac{1}{n}\sum_{i=1}\delta_{x_{i,c}}\mathbb{1}_{\{y_{i}=c\}}$. A discrepancy $\mathbb{F}(\mathbb{P})=D(\mathbb{P},\mathbb{Q})$ with $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ can be minimized by a Wasserstein over Wasserstein (WoW) gradient descent on this space. In particular, for $\mathbb{P}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}}$, the WoW gradient $\nabla_{\mathrm{W}_{\mathrm{W}_{2}}}\mathbb{F}(\mathbb{P})$ can be recovered by rescaling the Euclidean gradient of $F(\textbf{x})=\mathbb{F}(\mathbb{P})$ for $\textbf{x}:=(x_{i,c})_{i,c}$ by $nC$, i.e. $\nabla_{\mathrm{W}_{\mathrm{W}_{2}}}\mathbb{F}(\mathbb{P})(\mu_{c})(x_{i,c})=nC\nabla F(\textbf{x})$, see (Bonet et al., 2025a, Proposition B.7). Given $\mathbb{P}^{k}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}^{k}}$, $\mu_{c}^{k}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i,c}}$, the WoW gradient descent corresponds to updating each particle $x_{i,c}^{k}$ as
@@
We first focus on a toy example, where the target dataset contains $C=3$ classes, and each class forms a ring of $n=80$ samples. We learn a distribution $\mathbb{P}$ of the same form and with the same number of particles. On Figure 3, we show the convergence of the WoW gradient flows of $\mathrm{SOTDD}$, $\mathrm{SWB1DG}$ and $\mathrm{SWBG}$ with the 3 rings datasets as target, averaged over 100 different random batch of the target, and measured with the WoW distance. We use a step size of $\tau=1$ and 1000 iterations. The best performing distance appears to be $\mathrm{SWBG}$. We also report the particles along the flows of $\mathrm{SWBG}$ and $\mathrm{SOTDD}$ on Figure 2, clearly showing that the flow of $\mathrm{SWBG}$ converges faster than the one of $\mathrm{SOTDD}$.
@@
We also consider flowing image datasets to solve a $k$-shot transfer learning task (Alvarez-Melis and Fusi, 2021; Hua et al., 2023). In this task, we consider a target dataset $\mathbb{Q}$ with $k$ samples by class, where $k$ is typically small. To improve the classification results, we augment the dataset $\mathbb{Q}$ by concatenating it with samples flowed from another dataset $\mathbb{P}_{0}$, from which we have $n$ samples by class, still using the minimization of a distance on the dataset space. We focus here on $\mathbb{P}_{0}$ as the MNIST dataset (LeCun and Cortes, 2010), and $\mathbb{Q}$ as Fashion MNIST (Xiao et al., 2017) or USPS (Hull, 1994), and set $k\in\{1,5,10,100\}$ and $n=200$. Once the data has been flowed and each class of the flowed data has been matched to a class of the target data by solving the WoW OT problem, we train a LeNet5 on the augmented dataset. For the choice of the number of projections and iterations for the gradient descent, we perform a grid search. On Table 2, we report the best results of the trained classifier evaluated on a test set for OTDD, $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$, as well as the baseline where we only train on $\mathbb{Q}$. The results for OTDD and the baseline are taken from (Bonet et al., 2025a).
The results are overall comparable between $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$. Additional details, as well as examples of generated images, are provided in Appendix E.4. In particular, for a large number of iterations, both flows converge towards good images, but these do not necessarily provide the best results on the transfer task.
Table 2: Accuracy of a classifier on augmented datasets for $k\in\{1,10,10,100\}$. M refers to MNIST, F to Fashion MNIST and U to USPS.
@@ 6 CONCLUSION
We studied in this work in which cases the Busemann function is well defined on the Wasserstein space, and how to compute it in practice. Then, we showed that it can be used to define new efficient Sliced-Wasserstein distances on the space of datasets. Future works will include improving the scalability of $\mathrm{SWBG}$, e.g. by using Gaussian approximations with low-rank covariances (Bouveyron and Corneli, 2025), investigating other applications for the Busemann function such as Principal Component Analysis on the Wasserstein space (Cazelles et al., 2018; Vesseron et al., 2025), or its computation on probabilities over manifolds.
@@ Acknowledgements
CB thanks Pierre-Cyril Aubin-Frankowski for feedbacks on an earlier version of the draft. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015891 made by GENCI. CB and EC acknowledge the support of the Agence nationale de la recherche, through the PEPR PDE-AI project (ANR-23-PEIA-0004). NC was supported by the ANR chair OTTOPIA ANR-20-CHIA-0030.
@@ A.3 Backgound on the Wasserstein over Wasserstein Space
$\mathbb{F}(\mathbb{Q})=\mathbb{F}(\mathbb{P})+\iint\langle\xi(\mu)(x),y-x\rangle\ \mathrm{d}\gamma(x,y)\mathrm{d}\mathbb{\Gamma}(\gamma)+o\big(\mathrm{W}_{\mathrm{W}_{2}}(\mathbb{P},\mathbb{Q})\big).$ (42)
@@
To alleviate this computational burden, several approximations can be used. On one hand, it is possible to use a less costly distance as groundcost, such as the Sliced-Wasserstein distance [Baouan et al., 2025, Piening and Beinert, 2025a], reducing the complexity to $\mathcal{O}(C^{2}Ln\log n+C^{3}\log C)$. One could also use Linear OT [Wang et al., 2013, Liu et al., 2025] as groundcost, hence allowing to compute only $2C$ OT problems instead of $\mathcal{O}(C^{2})$ and reducing the complexity to $\mathcal{O}(Cn^{3}\log n)$. This is particularly appealing when there are lots of classes.
Piening and Beinert [2025b] very recently proposed a doubled slicing distance to compare distributions on $\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$. Given a projection $P^{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ for $\theta\in\Theta$, they first project in 1D the distributions $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ using $\varphi^{\theta}(\mu)=P^{\theta}_{\#}\mu\in\mathcal{P}_{2}(\mathbb{R})$, i.e. for $\mathbb{P}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, $\varphi^{\theta}_{\#}\mathbb{P}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R})\big)$. Then, they project $\mathbb{P}^{\theta}:=\varphi^{\theta}_{\#}\mathbb{P}$ in $\mathcal{P}_{2}(L^{2}([0,1])$ using, for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, $\phi(\mu)=F_{\mu}^{-1}$, and use the Sliced-Wasserstein distance on the Hilbert space $L^{2}([0,1])$ [Han, 2023]. More precisely, they define for $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$
@@
with the Sliced-Wasserstein distance defined on $\mathcal{P}_{2}\big(L^{2}([0,1])\big)$. For SW on this space, they use as slicing measure Gaussian on $\mathcal{P}_{2}\big(L^{2}([0,1])\big)$, which are sampled using Gaussian processes.

### v2

@@ 4.3 Slicing Datasets with Busemann
To leverage the closed-form between Gaussian (19), we use a Gaussian approximation of the classes, with a possible dimension reduction of the features beforehand as in (Hua et al., 2023). Let us denote $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$ a dimension reduction operator with $d^{\prime}\ll d$, and, for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d^{\prime}})$, the Gaussian approximation $\Xi(\mu)=\mathcal{N}\big(m(\mu),\Sigma(\mu)\big)$ with $m(\mu)=\int x\mathrm{d}\mu(x)$ and $\Sigma(\mu)=\int(x-m(\mu))\otimes(x-m(\mu))\ \mathrm{d}\mu(x)$ the mean and covariance operators. The label projections are then given by $Q^{\eta}(y)=B^{\eta}\big(\Xi\big(\psi_{\#}\varphi(y)\big)\big)$ with $\eta$ a geodesic ray on $\mathrm{BW}(\mathbb{R}^{d^{\prime}})$.
@@
with $\phi$ the cumulative distribution function of $\mathcal{N}(0,1)$. In practice, the geodesic ray can be sampled using $m_{1}\sim\mathcal{U}([-1,1])$ and setting $\sigma_{1}=\sqrt{1-m_{1}^{2}}$. Given $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, we define the Sliced-Wasserstein Busemann 1D Gaussian distance ($\mathrm{SWB1DG}$) as
@@
At the time of submission, we noticed the concurrent work (Piening and Beinert, 2026) whose sliced construction on $\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ is close to $\mathrm{SWB1DG}$, even though it is obtained from a different viewpoint. We detail in Appendix A.3 the relation between the two sliced distances.
@@ 5 EXPERIMENTS
Figure 2: Evolution of the WoW gradient flow of SWBG and SOTDD with the 3-rings dataset as target.
5.1 Correlation with OTDD
We report the results in Figure 1, where we first scatter the values obtained in ordinate for OTDD, and in abscissa for the sliced distances with 5000 projections. Then, we report the values of the Spearman and Pearson correlations.The Pearson correlation is equal to $\pm 1$ when both quantities are linearly correlated while Spearman correlation is equal to $\pm 1$ if the quantities are monotonically related. We observe that both $\mathrm{SWB1DG}$ and $\mathrm{SWBG}$ are better correlated to OTDD than $\mathrm{SOTDD}$. We note that the results do not match those in (Nguyen et al., 2025a), where they used only 10 pairs. We hypothesize that for such small number of samples, the Spearman and Pearson correlation are very sensitive to randomness. To verify the robustness of our results, we report in Table 1 the results obtained by bootstrapping 50 pairs of subdatasets, and averaging over 10 experiments. Using this process for different numbers of projections, we see that the Busemann based sliced distances always outperform $\mathrm{SOTDD}$, and often under a much smaller number of projections.
@@ 5.2 Flowing Labeled Datasets
Bonet et al. (2025a) modeled this task as a minimization problem over the space $\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, representing datasets with $n$ samples by class as $\mathbb{P}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ with $\mu_{c}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i,c}}\mathbb{1}_{\{y_{i}=c\}}$. A discrepancy $\mathbb{F}(\mathbb{P})=D(\mathbb{P},\mathbb{Q})$ with $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ can be minimized by a Wasserstein over Wasserstein (WoW) gradient descent on this space. In particular, for $\mathbb{P}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}}$, the WoW gradient $\nabla_{\mathrm{W}_{\mathrm{W}_{2}}}\mathbb{F}(\mathbb{P})$ can be recovered by rescaling the Euclidean gradient of $F(\textbf{x})=\mathbb{F}(\mathbb{P})$ for $\textbf{x}:=(x_{i,c})_{i,c}$ by $nC$, i.e. $\nabla_{\mathrm{W}_{\mathrm{W}_{2}}}\mathbb{F}(\mathbb{P})(\mu_{c})(x_{i,c})=nC\nabla F(\textbf{x})_{i,c}$, see (Bonet et al., 2025a, Proposition B.7). Given $\mathbb{P}^{k}=\frac{1}{C}\sum_{c=1}^{C}\delta_{\mu_{c}^{k}}$, $\mu_{c}^{k}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i,c}}$, the WoW gradient descent corresponds to updating each particle $x_{i,c}^{k}$ as
@@
Figure 3: Convergence of the flow towards the 3-rings, averaged over 100 random batches of the target.
@@
We first focus on a toy example, where the target dataset contains $C=3$ classes, and each class forms a ring of $n=80$ samples (Glaser et al., 2021; Bonet et al., 2025a). We learn a distribution $\mathbb{P}$ of the same form and with the same number of particles. In Figure 3, we show the convergence of the WoW gradient flows of $\mathrm{SOTDD}$, $\mathrm{SWB1DG}$ and $\mathrm{SWBG}$ with the 3-rings dataset as target, averaged over 100 different random batches of the target, and measured with the WoW distance (i.e. the OT problem with $\mathrm{W}_{2}^{2}$ as ground cost, which we denote $\mathrm{W}_{\mathrm{W}_{2}}$, see Appendix A.3). We use a step size of $\tau=1$ and 1000 iterations. The best performing distance appears to be $\mathrm{SWBG}$. We also report the particles along the flows of $\mathrm{SWBG}$ and $\mathrm{SOTDD}$ on Figure 2, clearly showing that the flow of $\mathrm{SWBG}$ converges faster than the one of $\mathrm{SOTDD}$.
@@
Table 2: Accuracy of a classifier on augmented datasets for $k\in\{1,5,10,100\}$. M refers to MNIST, F to Fashion MNIST and U to USPS.
@@
We now consider flowing image datasets to solve a $k$-shot transfer learning task (Alvarez-Melis and Fusi, 2021; Hua et al., 2023). In this experiment, we consider a target dataset $\mathbb{Q}$ with $k$ samples by class, where $k$ is typically small. To improve the classification results, we augment the dataset $\mathbb{Q}$ by concatenating it with samples flowed from another dataset $\mathbb{P}_{0}$, from which we have $n$ samples by class. This is done by minimizing a distance on the space of datasets with respect to $\mathbb{Q}$, i.e. by minimizing $\mathbb{F}(\mathbb{P})=\mathrm{D}(\mathbb{P},\mathbb{Q})$ starting from $\mathbb{P}_{0}$. Any divergence on $\mathcal{P}_{2}\big(\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})\big)$ may be chosen for $\mathrm{D}$. We focus here on comparing OTDD, SOTDD and SWB1DG for $\mathrm{D}$.
We take $\mathbb{P}_{0}$ as the MNIST dataset (LeCun and Cortes, 2010), and $\mathbb{Q}$ as Fashion MNIST (Xiao et al., 2017) or USPS (Hull, 1994), and set $k\in\{1,5,10,100\}$ and $n=200$. We minimize $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$ using the WoW gradient descent with step size $\tau=1$ and momentum $m=0.9$. For the choice of the number of projections for the Monte-Carlo approximation of the sliced distances, and of the number of iterations for the gradient descent, we perform a grid search. We refer to Appendix E.4 for more details about the experiment, and for the best parameters selected with the grid search.
Once the dataset $\mathbb{P}_{0}$ has been flowed towards $\mathbb{Q}$, and each class of the flowed dataset $\mathbb{P}_{T}$ has been matched to a class of the target dataset $\mathbb{Q}$ by solving the WoW OT problem, we train a LeNet5 on the augmented concatenated dataset. On Table 2, we report the accuracy of the classifier evaluated on a test set. We report the best results among the grid search for the classifier trained on $\mathbb{Q}$ augmented with data flowed by minimizing OTDD, $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$. We also report a baseline where the network was only trained on $\mathbb{Q}$. The results are averaged over 5 training of the networks, and 3 outputs of the flows for $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$, and are taken from (Bonet et al., 2025a) for OTDD and the baseline.
The results are overall comparable between $\mathrm{SWB1DG}$ and $\mathrm{SOTDD}$. Additional details, as well as examples of generated images, are provided in Appendix E.4. In particular, for a large number of iterations, both flows converge to good images. However, note that improved image quality does not necessarily translate into better performance on the transfer task.
Additionally, we report on Table 3 the runtimes for the transfer learning experiment, averaged over 3 outputs of the flows and trained for 5K epochs. We observe that SOTDD and SWB1DG almost have the same runtime, which is expected as they have the same computational complexity up to constants, and both are much more efficient than OTDD.
@@ 6 CONCLUSION
We studied in this work in which cases the Busemann function is well defined on the Wasserstein space, and how to compute it in practice. More precisely, we identified conditions to define geodesic rays on the Wasserstein space, showed that the Busemann function can be computed in general by solving an optimal transport problem, and derived closed-form formulas under which it can be computed efficiently. Then, we leveraged these closed-forms to define new efficient Sliced-Wasserstein distances on the space of datasets. Future works will include improving the scalability of $\mathrm{SWBG}$, e.g. by using Gaussian approximations with low-rank covariances (Bouveyron and Corneli, 2026), investigating other applications for the Busemann function such as Principal Component Analysis on the Wasserstein space (Cazelles et al., 2018; Vesseron et al., 2026), or its computation on probabilities over manifolds.
Table 3: Runtime in seconds for the transfer learning experiment from MNIST to Fashion MNIST.
Dataset $k$-shot OTDD SWB1DG SOTDD
M to F 1 $294.53\pm 5.21$ $13.53\pm 0.42$ $14.07\pm 0.37$
5 $1130.89\pm 108$ $13.84\pm 0.34$ $14.11\pm 0.21$
10 $2294.13\pm 48$ $14.00\pm 0.37$ $14.15\pm 0.18$
100 - $15.19\pm 0.49$ $15.31\pm 0.49$
@@ Acknowledgements
We thank the anonymous reviewers for their valuable comments. CB thanks Pierre-Cyril Aubin-Frankowski for feedbacks on an earlier version of the draft. This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015891 made by GENCI. CB and EC acknowledge the support of the Agence nationale de la recherche, through the PEPR PDE-AI project (ANR-23-PEIA-0004). NC was supported by the ANR chair OTTOPIA ANR-20-CHIA-0030 and contributes to AI Excellence Cluster SequoIA (grant ANR-23-IACL-0009).
@@ A.3 Backgound on the Wasserstein over Wasserstein Space
$\mathbb{F}(\mathbb{Q})=\mathbb{F}(\mathbb{P})+\iint\langle\xi(\pi^{1}_{\#}\gamma)(x),y-x\rangle\ \mathrm{d}\gamma(x,y)\mathrm{d}\mathbb{\Gamma}(\gamma)+o\big(\mathrm{W}_{\mathrm{W}_{2}}(\mathbb{P},\mathbb{Q})\big).$ (42)
@@
To alleviate this computational burden, several approximations can be used. On one hand, it is possible to use a less costly distance as groundcost, such as the Sliced-Wasserstein distance [Baouan et al., 2025, Piening and Beinert, 2025], reducing the complexity to $\mathcal{O}(C^{2}Ln\log n+C^{3}\log C)$. One could also use Linear OT [Wang et al., 2013, Liu et al., 2025] as groundcost, hence allowing to compute only $2C$ OT problems instead of $\mathcal{O}(C^{2})$ and reducing the complexity to $\mathcal{O}(Cn^{3}\log n)$. This is particularly appealing when there are lots of classes.
Piening and Beinert [2026] very recently proposed a doubled slicing distance to compare distributions on $\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$. Given a projection $P^{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ for $\theta\in\Theta$, they first project in 1D the distributions $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ using $\varphi^{\theta}(\mu)=P^{\theta}_{\#}\mu\in\mathcal{P}_{2}(\mathbb{R})$, i.e. for $\mathbb{P}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$, $\varphi^{\theta}_{\#}\mathbb{P}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R})\big)$. Then, they project $\mathbb{P}^{\theta}:=\varphi^{\theta}_{\#}\mathbb{P}$ in $\mathcal{P}_{2}(L^{2}([0,1])$ using, for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, $\phi(\mu)=F_{\mu}^{-1}$, and use the Sliced-Wasserstein distance on the Hilbert space $L^{2}([0,1])$ [Han, 2023]. More precisely, they define for $\mathbb{P},\mathbb{Q}\in\mathcal{P}_{2}\big(\mathcal{P}_{2}(\mathbb{R}^{d})\big)$
@@
with the Sliced-Wasserstein distance defined on $\mathcal{P}_{2}\big(L^{2}([0,1])\big)$. For SW on this space, they use Gaussian on $\mathcal{P}_{2}\big(L^{2}([0,1])\big)$ as slicing measure, which are sampled using Gaussian processes.